Question 1 (1 point) Given the following information, what is the location quotient for Mining and mineral extraction in Golden Eagle County? (Be sure to set your decimal places to four decimal places!) Total National Employment 111,970,095 National employment in mining and mineral extraction 581,582 Total employment in Golden Eagle County 56,000 Local employment in mining and mineral extraction 360 Your Answer
Answer #1

Location Quotient(LQ):

LQ is a term used in economic development.It is a simple ratio calculated to measure the concentration of a particular industry in a region in comparison to a larger region.it is used to compare an industries share of regional employment.

LQ is calculated as follows

LQi = ( ei / e ) / (Ei / E )

LQi = Location quotient for sector in the regional economy

ei = Employment in sector i in the regional economy.

e = Total employment in the local region

Ei = Employment in industry i in the national economy.

E = Total employment in the national economy.

The location quotient helps to calculate whether a region is employing enough workers in each sector as against a national economy.

If LQ=1, the employment is equal in the sector for the national and regional economy. So,the sector is non-basic and supply is equal to demand.

If LQ<1, the output is not sufficient to meet the local demand. It is also non-basic and imports are needed.

If LQ>1, output is more than sufficient to meet the local demand. It is basic and exporting the surplus is an option.

For the above problem LQ for mining and minerals extraction in golden eagle county is calculated as follows,

ei = 360

e = 56000

Ei = 581582

E = 111970095

LQi = (360/56000) / (581582/111970095)

= 0.006428 / 0.005194

= 1.23758

Here, LQi > 1 which means total employment in golden eagle county is more than sufficient to meet the local demand(mining and mineral extraction).